# Bibliography.bib

### v1

@@ 4. Isoperimetry and volume growth under $\mathrm{Ric}\geqslant 0$ and $\mathrm{biRic}\geqslant n-2$
    Now fix $\delta,V_{0}$. Let $o\in M$, let $K\subset M$ be a compact set such that $\mathrm{biRic}\geqslant n-2$ on $M\setminus K$, and take $r_{0}:=r_{0}(o,M,K)$ as in the statement of Lemma 6. Fix $r>r_{0}$ such that $\operatorname{vol}(B_{r}(o))>V_{0}$, and fix $\lambda<1$ such that $\lambda^{-(n-1)/2}\operatorname{vol}(\mathbb{S}^{n-1})<\operatorname{vol}(\mathbb{S}^{n-1})+\delta$.
    Applying Item 1 of Lemma 6 with the choice $\varepsilon=(n-2)(1-\lambda)$, we find a set $G$ with $E_{r+C\varepsilon^{-1}}\subset G\subset E_{r}$, such that $M\setminus G$ is bounded, and $\partial G$ is connected and satisfies $-\Delta_{\partial G}+\operatorname{Ric}_{\partial G}-(n-2)\lambda\geqslant 0$. When $n=4,5$, it follows from Theorem 1 (cf. (7)) that $\Omega:=M\setminus G$ satisfies
@@
    When $n=3$ (i.e., $\partial G$ is 2-dimensional), we may test the spectral condition $-\Delta_{\partial G}+\operatorname{Ric}_{\partial G}-\lambda\geqslant 0$ with the constant function $\psi\equiv 1$ and use Gauss–Bonnet theorem to get
@@
    Now, since $I$ is nondecreasing, if equality holds for some $v>0$ in (9), then it holds for every $V\geqslant v$. Thus, from [XingyuZhu, Theorem 3.2(3)] one gets that the quantity $\mathfrak{D}$ in there is equal to $\operatorname{vol}(\mathbb{S}^{n-1})$. Hence from the rigidity of [XingyuZhu, Theorem 3.2(2)] we get that there is a bounded open set $\Omega\subset M$ such that $M\setminus\Omega$ is isometric to $\partial\Omega\times[0,+\infty)$. Moreover, $\operatorname{vol}(\partial\Omega)=\operatorname{vol}(\mathbb{S}^{n-1})$ because isoperimetric regions with large volume exist and have boundary isometric to $\partial\Omega$. From the assumption $\mathrm{biRic}\geqslant n-2$ outside a compact set we get that $\mathrm{Ric}_{\partial\Omega}\geqslant n-2$. Thus, from the rigidity in the Bishop–Gromov comparison theorem, $\partial\Omega$ must be isometric to $\mathbb{S}^{n-1}$. The proof of Item 2 is thus concluded. ∎
@@
In the very recent paper [WeiXuZhang], the authors prove a sharp volume growth result at infinity when $n=3$ in Theorem 2. We stress that by joining Theorem 1, the Jacobian comparison argument in [WeiXuZhang, Lemma 5.3], and arguing as in [WeiXuZhang], one can prove the following generalization of [WeiXuZhang, Theorem 1.5]. Let $(M^{n},g)$ be a one-ended complete Riemannian manifold with $n\in\{3,4,5\}$, $\mathrm{Ric}_{M}\geqslant 0$, and $\mathrm{biRic}_{M}\geqslant n-2$ outside a compact set. Then, for any $o\in M$ it holds
@@
We also present here a proof of (57) using a different limit space argument. In this argument, we need to assume the noncollapsing condition $\inf_{p\in M}\operatorname{vol}(B_{1}(p))>0$.
Take a radial geodesic $\gamma$ issuing from $o$, and let $b$ be the associated Busemann function. Since $M$ has linear volume growth by Item 1 in Theorem 2, and since $\mathrm{Ric}\geqslant 0$, we can use [Sormani, Corollary 23] and infer $0\geqslant m:=\inf_{p\in M}b(p)>-\infty$. In addition, by [XingyuZhu, Proposition 2.6], we get that there is a constant $C>0$ such that $d(p,\gamma)\leqslant C$ for every $p\in M$. Hence we can freely apply [XingyuZhu, Proposition 2.10], and [XingyuZhu, Theorem 3.2] in what follows.

### v2

@@ 4. Isoperimetry under $\mathrm{Ric}\geqslant 0$ and spectral $\mathrm{biRic}\geqslant n-2$
    Now fix $\delta,V_{0}$. Let $o\in M$, let $K\subset M$ be a compact set such that $-\gamma\Delta+\mathrm{biRic}\geqslant n-2$ on $M\setminus K$, and take $r_{0}:=r_{0}(o,M,K)$ as in the statement of Lemma 6. Fix $r>r_{0}$ such that $\operatorname{vol}(B_{r}(o))>V_{0}$, and fix $\lambda<1$ such that $\lambda^{-(n-1)/2}\operatorname{vol}(\mathbb{S}^{n-1})<\operatorname{vol}(\mathbb{S}^{n-1})+\delta$.
    Applying Item 1 of Lemma 6 with the choice $\varepsilon=(n-2)(1-\lambda)$, we find a set $G$ with $E_{r+C\varepsilon^{-1}}\subset G\subset E_{r}$, such that $M\setminus G$ is bounded, and $\partial G$ is connected and satisfies $-\frac{4}{4-\gamma}\Delta_{\partial G}+\operatorname{Ric}_{\partial G}-(n-2)\lambda\geqslant 0$. When $n=4,5$, it follows from Theorem 1 (cf. (55)) that $\Omega:=M\setminus G$ satisfies
@@
    When $n=3$ (i.e., $\partial G$ is 2-dimensional), we may test the spectral condition $-\frac{4}{4-\gamma}\Delta_{\partial G}+\operatorname{Ric}_{\partial G}-\lambda\geqslant 0$ with the constant function $\psi\equiv 1$ and use Gauss–Bonnet theorem to get
@@
    Now, since $I$ is nondecreasing, if equality holds for some $v>0$ in (7), then it holds for every $V\geqslant v$. Thus, from [XingyuZhu, Theorem 3.2(3)] one gets that the quantity $\mathfrak{D}$ in there is equal to $\operatorname{vol}(\mathbb{S}^{n-1})$. Hence from the rigidity of [XingyuZhu, Theorem 3.2(2)] we get that there is a bounded open set $\Omega\subset M$ such that $M\setminus\Omega$ is isometric to $\partial\Omega\times[0,+\infty)$. Moreover, $\operatorname{vol}(\partial\Omega)=\operatorname{vol}(\mathbb{S}^{n-1})$ because isoperimetric regions with large volume exist and have boundary isometric to $\partial\Omega$. From the assumption $-\gamma\Delta+\mathrm{biRic}\geqslant n-2$ outside a compact set we get that $-\gamma\Delta_{\partial\Omega}+\mathrm{Ric}_{\partial\Omega}\geqslant n-2$. Thus, from the rigidity in Theorem 1, see in particular Remark 5, $\partial\Omega$ must be isometric to $\mathbb{S}^{n-1}$. The proof of Item 2 is thus concluded. ∎
@@
In the papers [WeiXuZhang, HuangLiu], which appeared when we were completing a first draft of this paper, the authors prove a sharp volume growth result at infinity when $n=3$, and $\gamma=0$ in Theorem 2.
Joining our result in Theorem 2, and using a different limit space argument, we can prove the following. Let $(M^{n},g)$ be a one-ended complete Riemannian manifold with $n\in\{3,4,5\}$, $0\leqslant\gamma<6-n$ (or $0\leqslant\gamma\leqslant 2$ when $n=4$), $\mathrm{Ric}\geqslant 0$, and $\lambda_{1}(-\gamma\Delta+\mathrm{biRic})\geqslant n-2$ outside a compact set. Assume further
$\inf_{p\in M}\operatorname{vol}(B_{1}(p))>0.$ (79)
Then, for any $o\in M$ it holds
@@
Note that in dimension $n=3$, and when $\gamma=0$, (79) follows from the other assumptions, see the recent [HuangLiu, Proposition 1.10]. In general, it is likely that the assumption (79) is not needed: e.g., by joining Theorem 1, the Jacobian comparison argument in [WeiXuZhang, Lemma 5.3], and arguing as in [WeiXuZhang]; or by proving a more general version of [HuangLiu, Proposition 1.10].
Let us present a proof of (80) under the non-collapsing assumption (79). With (79) available, we are free to use the results of [XingyuZhu]. Take a radial geodesic $\gamma$ issuing from $o$, and let $b$ be the associated Busemann function. Since $M$ has linear volume growth by Item 1 in Theorem 2, and since $\mathrm{Ric}\geqslant 0$, we can use [Sormani, Corollary 23] and infer $0\geqslant m:=\inf_{p\in M}b(p)>-\infty$. In addition, by [XingyuZhu, Proposition 2.6], we get that there is a constant $C>0$ such that $d(p,\gamma)\leqslant C$ for every $p\in M$. Hence we can freely apply [XingyuZhu, Proposition 2.10], and [XingyuZhu, Theorem 3.2] in what follows.